Hi there
I have a curly trigonometry question that I simply do not understand as it is written in such a way as I have never seen before. Any assistance would be greatly appreciated. Thank you.
It asks me to evaluate tan^2 2pie/3.
Thank you so very much.

Answer by stanbon(48546)

evaluate tan^2 2pie/3
tan[(2/3)pi] = tan 120 degrees = -1.73205... or -sqrt(3)
-------
[tan[(2/3)pi]^2 = (-1.73205...)^2 = 3
=================
Cheers,
Stan H.